Given: x = 1 – cost and y = t – sint. Find y’.


1
Expert's answer
2022-02-23T08:23:16-0500
"y'=\frac{dy}{dx}=\frac{dy/dt}{dx/dt}=\frac{1-\cos t}{\sin t}"